# GeoRecon: Graph-Level Representation Learning for 3D Molecules via Reconstruction-Based Pretraining

### v1

@@ 3.1 Model Architecture
Shared Equivariant Encoder At the core of GeoRecon is an 8-layer SE(3)-equivariant transformer encoder adapted from TorchMD-Net (Thölke and De Fabritiis, 2022). The encoder operates on molecular graphs defined by atomic numbers $\{z_{i}\}$ and spatial coordinates $\{\bm{r}_{i}^{*}\}$.
Each node is initialized via a learnable embedding of its atomic type, optionally enriched through distance-aware neighbor aggregation. Edge features are computed by applying radial basis function (RBF) expansion to interatomic distances, modulated by a smooth cosine cutoff, while directional geometry is encoded by the normalized displacement vector:
$\bm{f}_{ij}=\text{{RBF}}(\|\bm{\mathrm{d}}_{ij}\|)\cdot\text{{Cutoff}}(\|\bm{\mathrm{d}}_{ij}\|).$
Figure 1: Overview of the GeoRecon framework. Given a molecular structure with atom types and 3D coordinates, the model encodes it using eight layers of SE(3)-equivariant attention. A pooled graph-level representation is fused with node embeddings and passed to a lightweight decoder to predict scaled noise. Additional denoising and clean heads regularize local features during pretraining. The pretrained encoder is then finetuned for downstream molecular property prediction tasks.
At each layer, scalar and vector-valued features are jointly updated via attention-weighted message passing over neighboring nodes, incorporating edge attributes and directional information. The entire layer can be compactly written as:
$(\bm{h}_{i}^{(l+1)},$ $\ \bm{u}_{i}^{(l+1)})=\textsc{EquiAttn}^{(l)}\left(\bm{h}_{i}^{(l)},\right.$
$\left.\ \bm{u}_{i}^{(l)},\ \{\bm{h}_{j}^{(l)},\ \bm{f}_{ij},\ \hat{\bm{d}}_{ij}\}_{j\in\mathcal{N}(i)}\right),$
where $\bm{h}_{i}^{(l)}\in\mathbb{R}^{d}$ and $\bm{u}_{i}^{(l)}\in\mathbb{R}^{3\times d}$ denote the scalar and vector representations, respectively. For simplicity, we omit the task-specific superscript ()(*)( ∗ ) in the notations, as all tasks are treated in the same way. Residual connections and layer normalization are applied at each layer.
Denoising Head The denoising head is designed to reverse the synthetic perturbation applied during pretraining. Specifically, random Gaussian noise $\bm{\epsilon}$ is added to the 3D coordinates of each atom at the data processing stage. The model is trained to predict this noise vector at each node with a mean squared error (MSE) loss:
$\hat{\bm{\epsilon}}_{i}=\mathrm{MLP}_{\texttt{denoise}}(\bm{h}_{i})$
$\quad\mathcal{L}_{\texttt{denoise}}=\frac{1}{N}\sum_{i}^{N}\|\hat{\bm{\epsilon}}_{i}-\bm{\epsilon}_{i}\|_{2}^{2}\sim\frac{\sigma^{2}}{k_{B}T}\bm{F}$
According to Equation (1), the equivalence between coordinate denoising and force field learning under the Boltzmann distribution assumption can be readily established. A brief derivation is included in the appendix A for completeness and reference. For clarity, the symbol $\sim$ above denotes that the denoising objective form is approximately equivalent to learning the molecular force field, up to a constant factor. That is, minimizing the MSE between predicted and true noise vectors corresponds to estimating the gradient of the potential energy surface, which governs atomic forces in physical systems. This interpretation is also supported by the general connection between denoising and score matching established in Vincent (2011).
@@
To fully leverage the clean molecular conformations already processed during pretraining, we integrate them into the denoising pipeline as well. Specifically, clean molecules—with zero perturbation—are passed through the same encoder and denoising head. Their supervision target is a zero noise vector $\bm{\epsilon}_{i}=\bm{0}$, naturally enforcing representational consistency between clean and noised structures. This unifies the treatment of clean and noised molecules under a single objective, regularizing the encoder without requiring additional components or losses. The denoising loss on clean structures is denoted by $\mathcal{L}_{\texttt{clean}}.$ To evaluate the impact of the auxiliary clean alignment task, we conduct additional ablation studies in Appendix 6.
Reconstruction Head with Graph-Level Conditioning The reconstruction head in GeoRecon is designed to predict the coordinate noise applied to heavily perturbed molecular inputs, rather than directly reconstructing the original coordinates. It leverages global semantic information $\bm{g}$ distilled from clean conformations to guide this denoising process. Unlike standard autoencoders that decode absolute positions, our approach treats reconstruction as a tougher denoising task—applying stronger noise to the input and predicting the corresponding perturbation under the guidance of a lightweight decoder conditioned on the graph-level embedding. This formulation not only avoids ambiguity due to global orientation, but also promotes the learning of structure-aware representations aligned with whole-molecule geometry.
To construct the conditioning signal, the clean input undergoes a complete forward pass through the encoder, producing node-level representations $\{\bm{h}_{i}^{\texttt{(clean)}}\}$. These are aggregated via mean pooling to yield a graph-level vector:
$\bm{g}:=\textsc{Aggr}\left(\bigoplus_{i=1}^{N}\bm{h}_{i}^{\texttt{(clean)}}\right)=\frac{1}{N}\sum_{i=1}^{N}\bm{h}_{i}^{\texttt{(clean)}}.$
@@
The resulting vector $\bm{g}$ encapsulates holistic structural semantics and is broadcasted to all nodes in the reconstruction input. We adopt mean pooling for its simplicity and stability under coordinate perturbations, though more sophisticated pooling strategies are compatible with our framework. Each node concatenates this global vector with its local features $\bm{h}_{i}^{(\texttt{rec})}$, forming the input to a lightweight decoder:
$\bm{z}_{i}=\textsc{Concat}(\bm{g},\bm{h}_{i}^{\texttt{(rec)}}),\quad\hat{\bm{\epsilon}}_{i}^{\texttt{(rec)}}=\textsc{MLP}_{\texttt{rec}}(\bm{z}_{i}),\quad\mathcal{L}_{\texttt{rec}}=\frac{1}{N}\sum_{i}^{N}\|\hat{\bm{\epsilon}}_{i}^{\texttt{(rec)}}-\lambda\cdot\bm{\epsilon}_{i}\|_{2}^{2}.$
Here, $\lambda$ is the fixed scaling factor that determines the difficulty of the reconstruction task, as introduced above. By increasing the noise level in the reconstruction targets $\bm{r}_{i}^{\texttt{rec}}$, the model is encouraged to place greater emphasis on global context during decoding. This design implicitly guides the encoder to embed longer-range geometric dependencies into the graph-level representation $\bm{g}$, enhancing its ability to capture coherent molecular structures beyond local neighborhoods.
@@
As a result, the learned latent space becomes both locally descriptive—sensitive to fine-grained atomic perturbations—and globally coherent, reflecting the overall molecular topology. This dual expressiveness contributes to improved generalization in downstream molecular property prediction tasks, particularly those sensitive to spatial organization, such as total energy, dipole moment, and orbital energy estimation.
@@ 3.2 Multi-Task Training Objective
$\mathcal{L}_{\texttt{total}}=\lambda_{\texttt{denoise}}\mathcal{L}_{\texttt{denoise}}+\lambda_{\texttt{rec}}\mathcal{L}_{\texttt{rec}}+\lambda_{\texttt{clean}}\mathcal{L}_{\texttt{clean}}.$
Each component provides complementary supervision: the denoising loss captures local geometric regularities, the reconstruction loss promotes global structural awareness, and the clean alignment term is introduced to reuse the clean conformations during pretraining. Together, these tasks encourage the encoder to learn transferable representations that are effective across both node-level and graph-level downstream objectives. The loss weights are empirically chosen to ensure that the magnitudes of individual loss terms are balanced during training.
@@ 4 Experiments
4.1 Basic Setup
Training Settings. We pretrain GeoRecon on a large-scale 3D molecular datasets PCQM4Mv2 (Hu et al., 2021). Pretraining uses the Adam optimizer with a learning rate of $4\times 10^{-4}$ and cosine decay scheduling over 400k steps, with a batch size of 50. The encoder consists of 8 SE(3)-equivariant attention layers with 256-dimensional embeddings and 8 attention heads. During pretraining, Gaussian noise with a standard deviation of 0.04 Å is added to atomic coordinates. During finetuning, all tasks share the same pretrained encoder, followed by task-specific heads and loss functions adapted to the target property type. We adopt a full fine-tuning protocol where all model parameters are updated during downstream training.
All main experiments were conducted on NVIDIA RTX 4090 GPUs. Specifically, pretraining was performed using 4× RTX 4090 GPUs (24GB each) over 400k steps, while downstream fine-tuning on each task was carried out using a single RTX 4090 GPU.
@@ 4.1 Basic Setup
Baselines. Our method is built upon the denoising task, with an additional reconstruction auxiliary task. Therefore, our primary direct comparison is with the denoising baseline Coord (Zaidi et al., 2023). To clearly highlight improvements over this baseline, we mark them in green in Tables 1 and 2. Furthermore, we include comparisons to traditional methods without pre-training to demonstrate the benefits of pretraining, as well as to several pre-training approaches:
  • Training from scratch: SchNet (Schütt et al., 2018) employs continuous-filter convolutions to model quantum interactions in molecules. EGNN (Satorras et al., 2021) introduces equivariant message passing under Euclidean transformations. DimeNet (Gasteiger et al., 2020b) incorporate directional information via angular message passing. DimeNet++ (Gasteiger et al., 2020a) is an enhanced version of DimeNet, offering improved efficiency and accuracy. PaiNN (Schütt et al., 2021) utilizes equivariant message passing tailored for predicting tensorial properties and molecular spectra. SphereNet (Liu et al., 2021b) enhances spatial reasoning using spherical coordinates. TorchMD-Net (Thölke and De Fabritiis, 2022) applies equivariant transformers to molecular dynamics tasks.
  • Pretraining then finetune: Transformer-M (Luo et al., 2022) is a versatile model capable of processing both 2D and 3D molecular data through separate encoding channels. SE(3)-DDM (Liu et al., 2022) pretrains via denoising distance matching under SE(3)-invariance. 3D-EMGPJiao et al. (2023) leverages energy-motivated objectives for equivariant pretraining. Coord (Zaidi et al., 2023) applies coordinate denoising for geometry-aware representation learning. MolSpectra (Wang et al., 2025) introduces spectral prediction as an auxiliary task, enriching molecular representations with quantum-informed supervision.
Although MolSpectra also builds upon Coord codebase, it incorporates auxiliary spectral information derived from analytical force fields during pretraining. This reliance on external physical supervision makes it methodologically distinct from our purely geometry-driven and task-agnostic framework, and thus not suitable as a core baseline for our evaluation.
4.2 QM9: Prediction of Various Molecular Properties
We first evaluate our method on QM9 dataset. QM9 (Ramakrishnan et al., 2014) is a widely-used benchmark for quantum machine learning on small organic molecules, containing $\sim$134k samples composed of H, C, N, O, F atoms. Each sample provides an equilibrium 3D geometry along with 12 quantum chemical properties computed via DFT. These include dipole moment ($\mu$), isotropic polarizability ($\alpha$), frontier orbital energies (HOMO, LUMO, and their gap), enthalpies ($U,U_{0},H,G$), zero-point vibrational energy (ZPVE), electronic spatial extent ($R^{2}$), and heat capacity at 298.15K ($C_{v}$), all treated as regression targets. We follow the standard data split with 110k molecules for training, 10k for validation, and the remaining for testing. Mean Absolute Error (MAE) is used as the evaluation metric. For each target property, we attach a scalar regression head to the pretrained encoder and fine-tune the entire model.
See Table 1 for results. As shown, GeoRecon achieves strong performance across a range of molecular property prediction tasks. GeoRecon achieves the best performance across all baselines on the $U$, $U_{0}$, and $H$ tasks. Compared to our backbone Coord, GeoRecon shows noticeable improvements on key targets including $\mu$, $\alpha$, HOMO, LUMO, and gap. Interestingly, it also outperforms MolSpectra, the best performing baseline, on ZPVE $R^{2}$, despite not relying on external spectral data. These results support the effectiveness of incorporating graph-level reconstruction in pretraining and demonstrate that our method enhances the encoder’s ability to capture both fine-grained and global molecular structure.
@@ 4.2 QM9: Prediction of Various Molecular Properties
Table 1: MAE (↓) on QM9 property prediction. Best and second-best results are bolded and underlined, respectively. Aver-rank is the mean of per-metric ranks. The methods are divided into two groups: training from scratch and pre-training then fine-tuning. Green text indicates the relative improvement (%) over the base model; red indicates degradation. During pretraining, the loss weights $\lambda_{\texttt{denoise}}$, $\lambda_{\texttt{rec}}$, and $\lambda_{\texttt{clean}}$ were set to 1, 0.5, and 10, respectively.
$\mu$ $\alpha$ homo lumo gap $R^{2}$ ZPVE $U_{0}$ $U$ $H$ $G$ $C_{v}$ Average
(D) ($a_{0}^{3}$) (meV) (meV) (meV) ($a_{0}^{2}$) (meV) (meV) (meV) (meV) (meV) ($\tfrac{\text{cal}}{\text{mol}\cdot\text{K}}$) rank
SchNet 0.033 0.235 41.0 34.0 63.0 0.070 1.70 14.00 19.00 14.00 14.00 0.033 10.83
EGNN 0.029 0.071 29.0 25.0 48.0 0.106 1.55 11.00 12.00 12.00 12.00 0.031 10.16
DimeNet++ 0.030 0.044 24.6 19.5 32.6 0.330 1.21 6.32 6.28 6.53 7.56 0.023 6.33
PaiNN 0.016 0.045 27.6 20.4 45.7 0.070 1.28 5.85 5.83 5.98 7.35 0.024 5.33
@@
TorchMD-Net 0.011 0.059 20.3 17.5 36.1 0.033 1.84 6.15 6.38 6.16 7.62 0.026 5.92
Transformer-M 0.037 0.041 17.5 16.2 27.4 0.075 1.18 9.37 9.41 9.39 9.63 0.022 5.92
SE(3)-DDM 0.015 0.046 23.5 19.5 40.2 0.122 1.31 6.92 6.99 7.09 7.65 0.024 7.00
3D-EMGP 0.020 0.057 21.3 18.2 37.1 0.092 1.38 8.60 8.60 8.70 9.30 0.026 7.67
MolSpectra 0.011 0.048 15.5 13.1 26.8 0.410 1.71 5.67 5.45 5.87 6.18 0.021 3.33
Coord 0.016 0.052 17.7 14.7 31.8 0.450 1.71 6.57 6.11 6.45 6.91 0.020 5.67
GeoRecon (Ours) 0.012 $\downarrow$25.000.050 $\downarrow$3.8516.6 $\downarrow$6.2113.7 $\downarrow$6.8030.6 $\downarrow$3.770.250 $\downarrow$44.441.42 $\downarrow$16.965.64 $\downarrow$14.165.36 $\downarrow$12.275.55 $\downarrow$13.956.79 $\downarrow$1.740.023 $\uparrow$15.003.50
*: denotes using external data
4.3 MD17: Molecular Dynamic Prediction
We then evaluate the model on MD17 (Chmiela et al., 2017), which is a benchmark dataset for learning energy and force fields from molecular dynamics (MD) simulations. It contains ab initio molecular dynamics trajectories for small organic molecules, with atom-wise force labels and associated energy values provided for each conformation. Unlike QM9, which focuses on equilibrium properties, MD17 emphasizes modeling local geometric sensitivity under thermal perturbations, making it a challenging testbed for fine-grained prediction. We evaluate on eight commonly used molecules. For each molecule, we adopt the standard low-data regime—training on 1,000 samples, validating on 50, and testing on the remaining configurations. Following prior work, we use Mean Absolute Error (MAE) on the predicted atomic forces as the evaluation metric.
As shown in Table 2, GeoRecon demonstrates strong performance on the MD17 benchmark for force prediction. It shows the best performance on multiple molecules, including Aspirin (0.068), Benzene (0.048), and Salicylic Acid (0.069), and ranks among the top three on others such as Ethanol and Malonaldehyde. While not consistently the best across all targets, our method performs competitively against both pretraining-based and fully supervised baselines. These results support the effectiveness of incorporating graph-level reconstruction in modeling fine-grained molecular dynamics, and suggest that integrating global structural information can enhance local force prediction accuracy without additional overhead.
Table 2: MAE (↓) of force prediction on MD17 dataset (kcal/mol/Å). Best and second-best results are bolded and underlined, respectively. Aver-rank is the mean of per-metric ranks. The methods are divided into two groups: training from scratch and pre-training then fine-tuning. Green text indicates the relative improvement (%) over the base model; red indicates degradation. During pretraining, the loss weights $\lambda_{\texttt{denoise}}$, $\lambda_{\texttt{rec}}$, and $\lambda_{\texttt{clean}}$ were set to 1, 0.5, and 10, respectively.
Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic Acid Toluene Uracil Aver-rank
SphereNet 0.430 0.178 0.208 0.340 0.178 0.360 0.155 0.267 6.38
SchNet 1.350 0.310 0.390 0.660 0.580 0.850 0.570 0.560 8.75
DimeNet 0.499 0.187 0.230 0.383 0.215 0.374 0.216 0.301 7.63
PaiNN 0.338 0.224 0.319 0.077 0.195 0.094 0.139 -
TorchMD-Net 0.245 0.219 0.107 0.167 0.059 0.128 0.064 0.089 3.50
SE(3)-DDM 0.453 0.166 0.288 0.129 0.266 0.122 0.183 -
MolSpectra 0.099 0.097 0.052 0.077 0.085 0.094 0.075 0.095 2.50
Coord 0.211 0.169 0.096 0.139 0.053 0.109 0.058 0.074 2.25
GeoRecon (Ours) 0.068 $\downarrow$67.770.048 $\downarrow$71.60.057 $\downarrow$40.630.093 $\downarrow$33.090.073 $\uparrow$37.740.069 $\downarrow$36.70.068 $\uparrow$17.240.098 $\uparrow$32.432.13
*: denotes using external data
4.4 Ablation Study

### v2

@@ 4.2 Model Architecture
Shared Equivariant Encoder At the core of GeoRecon is an 8-layer SE(3)-equivariant transformer encoder adapted from TorchMD-Net (Thölke & De Fabritiis, 2022). The encoder operates on molecular graphs defined by atomic numbers $\{z_{i}\}$ and spatial coordinates $\{\bm{r}_{i}^{*}\}$.
Denoising Head The denoising head is designed to reverse the synthetic perturbation applied during pretraining. We deliberately follow the node-level denoising paradigm to keep consistency to previous work and decrease migration difficulty, but our objectives explicitly enforce utilizing global graph structure.
According to Theorem 1, the equivalence between coordinate denoising and force field learning under the Boltzmann distribution assumption can be readily established. A brief derivation is included in the appendix D for completeness and reference.
@@
To fully leverage the clean molecular conformations already processed during pretraining, we integrate them into the denoising pipeline as well. Specifically, clean molecules, with zero perturbation, are passed through the same encoder and denoising head. Their supervision target is a zero noise vector, naturally enforcing representational consistency between clean and noised structures. This unifies the treatment of clean and noised molecules under a single objective, regularizing the encoder without requiring additional components or losses. To evaluate the impact of the auxiliary clean alignment task, we conduct additional ablation studies in Appendix E.
Reconstruction Head with Graph-Level Conditioning To mitigate the ambiguity introduced by global orientation, we design the reconstruction head differently from standard autoencoders. Instead of directly decoding absolute coordinates, GeoRecon formulates reconstruction as a more challenging denoising task: stronger noise is applied to the inputs, and the model predicts the corresponding perturbation with a lightweight decoder conditioned on the graph-level embedding. This design encourages the learning of structure-aware representations aligned with whole-molecule geometry, while leveraging global semantic information $\bm{g}$ derived from clean conformations to guide the denoising process. Intuitively, as the perturbation is very strong, the node embeddings after the encoder become extremely noisy. This time, if we directly concatenate the clean node embeddings with them, the decoder is easy to recover the added noise. Yet, we choose to not feed the clean node embeddings, but use the pooled clean graph embedding instead. This enforces the graph embedding to contain rich local information and be coherent with node-level representations. We make the decoder lightweight to further increase the task difficulty and thus reinforcing the reliance on the graph-level embeddings.
@@
To construct the conditioning signal, the clean input undergoes a complete forward pass through the encoder, producing node-level representations $\{\bm{h}_{i}^{\texttt{(cln)}}\}$. These are aggregated via pooling function to yield a graph-level vector. The resulting vector $\bm{g}$ encapsulates holistic structural semantics and is broadcasted to all nodes in the reconstruction input. We adopt mean pooling for its simplicity and stability under coordinate perturbations, while noting that more sophisticated pooling strategies are also compatible with our framework. Each node concatenates this global vector with its local representation $\bm{h}_{i}^{(\texttt{rec})}$, forming the input to a lightweight decoder:
$\bm{z}_{i}=\textsc{Concat}(\bm{g},\bm{h}_{i}^{\texttt{(rec)}}),\quad\hat{\bm{\epsilon}}_{i}^{\texttt{(rec)}}=\textsc{MLP}_{\texttt{rec}}(\bm{z}_{i}).$
@@
Consequently, the learned latent space attains dual expressiveness: it remains locally descriptive, capturing fine-grained atomic perturbations, while also being globally coherent in reflecting the overall molecular topology. This property improves generalization in downstream molecular property prediction, especially for tasks that depend strongly on spatial organization, including total energy, dipole moment, and orbital energy estimation.
@@ 4.3 Multi-Task Training Objective
$\mathcal{L}_{\texttt{total}}=\lambda_{\texttt{nsd}}\mathcal{L}_{\texttt{nsd}}+\lambda_{\texttt{rec}}\mathcal{L}_{\texttt{rec}}+\lambda_{\texttt{cln}}\mathcal{L}_{\texttt{cln}}.$
Each loss term provides complementary supervision, encouraging the encoder to learn representations that are both locally descriptive and globally coherent. The loss weights are empirically chosen to balance the magnitudes of individual terms during training.
@@ 5 Experiments
Information about training devices, hyperparameters, and computational costs are provided in Appendix J. Training code is available in the supplementary materials.
Figure 2: Representation stability of GeoRecon (upper row) vs. Coord (lower row) near equilibrium conformations ($\|\delta\bm{x}\|\leq 1$ Å). Left: averages over multiple molecules. Middle & Right: 2D perturbation heatmaps for a randomly selected PCQM4Mv2 sample, where the horizontal and vertical axes correspond to the magnitudes of two coordinate perturbations applied to a random atom, and the color encodes the norm change of the representation ($\|\delta\bm{h}\|$). Larger blue regions indicate higher stability of the representation near the equilibrium conformation. The red curve in the middle column highlights $\delta y=0$ ($\|\delta\bm{h}\|$ along the $x$-axis). Scale bars: shared on the left (left column) and at the bottom (middle).
5.1 Lipshitz Constant Analyze after Pretraining
We measured the Lipshitz constant of GeoRecon after pretraining on PCQM4Mv2 and observed a substantial reduction ($\sim$ 99%) in GeoRecon compared to Coord.
As shown in Table 1 , GeoRecon consistently achieves median $L(x)\approx$ 30 with negligible variance across step sizes, while the Coord baseline yields values on the order of $10^{4}\sim 10^{5}$, representing a reduction of roughly three orders of magnitude in the local Lipschitz constant.
Figure 2 compares the sensitivity of GeoRecon and Coord encoders near equilibrium conformations. Under small perturbations ($\|\delta\bm{x}\|\leq 0.4$ Å), GeoRecon exhibits much smoother representation manifolds, with smaller changes in $\|\bm{h}\|$ compared to Coord. This trend is consistent both in individual samples and in averages across multiple molecules.
The experimental results confirm our intuition that anchoring a graph-level perspective during the pre-training stage leads to a smoother latent space. This provides a stronger theoretical foundation for our improvements in downstream tasks, as smoother representation spaces have been shown to facilitate better learning procedures in various other settings (Lee et al., 2025; Guo et al., 2024; Zhang et al., 2025; Krishnan et al., 2020).
5.2 Test on Downstream Tasks
@@
Our method is built upon the denoising task, with an additional reconstruction auxiliary task. Therefore, our primary direct comparison is with the denoising baseline Coord (Zaidi et al., 2023). To clearly highlight improvements over this baseline, we mark them in green in Tables 2,  3 and 4.
We include comparisons to traditional methods without pre-training to demonstrate the benefits of pretraining, as well as to several pre-training approaches.
@@
Table 2: MAE (↓) on QM9 property prediction. Best and second-best results are bolded and underlined, respectively. Aver-rank is the mean of per-metric ranks. The methods are divided into two groups: training from scratch and pre-training then fine-tuning. Green text indicates the relative improvement (%) over the base model; red indicates degradation.
Task $\mu$ $\alpha$ homo lumo gap $R^{2}$ ZPVE $_{0} UU$ $H$ $G$ $C_{v}$ Average
Unit (D) ($a_{0}^{3}$) (meV) (meV) (meV) ($a_{0}^{2}$) (meV) (meV) (meV) (meV) (meV) $\left(\tfrac{\text{cal}}{\text{mol}\cdot\text{K}}\right)$ rank
SchNet 0.033 0.235 41.0 34.0 63.0 0.070 1.70 14.00 19.00 14.00 14.00 0.033 9.33
EGNN 0.029 0.071 29.0 25.0 48.0 0.106 1.55 11.00 12.00 12.00 12.00 0.031 9.00
DimeNet++ 0.030 0.044 24.6 19.5 32.6 0.330 1.21 6.32 6.28 6.53 7.56 0.023 6.84
PaiNN 0.012 0.045 27.6 20.4 45.7 0.070 1.28 5.85 5.83 5.98 7.35 0.024 5.67
@@
TorchMD-Net 0.011 0.059 20.3 17.5 36.1 0.033 1.84 6.15 6.38 6.16 7.62 0.026 4.17
Transformer-M 0.037 0.041 17.5 16.2 27.4 0.075 1.18 9.37 9.41 9.39 9.63 0.022 3.67
SE(3)-DDM 0.015 0.046 23.5 19.5 40.2 0.122 1.31 6.92 6.99 7.09 7.65 0.024 6.33
3D-EMGP 0.020 0.057 21.3 18.2 37.1 0.092 1.38 8.60 8.60 8.70 9.30 0.026 6.00
[…]
Relative Gain 23.4% 7.0% 4.8% 3.5% 1.3% 47.2% 11.9% 21.7% 16.7% 20.5% 7.5% -10.0%
5.2.1 Evaluation on QM9
QM9 benchmark (Ramakrishnan et al., 2014) contains 134k stable small organic molecules (up to 9 heavy atoms: C, O, N, F, and H), each optimized at the B3LYP/6-31G(2df,p) level of DFT. It provides geometric, energetic, electronic, and thermodynamic properties; in this work, we focus on the 3D geometries and associated properties relevant for molecular representation learning.
See Table 2 for results. The details of baselines are shown in Appendix I. As shown, GeoRecon achieves strong performance across a range of molecular property prediction tasks. It achieves the best performance across all baselines on multiple targets, including $U$, $U_{0}$, $H$, and $G$. Compared to our backbone Coord, GeoRecon also shows noticeable improvements on several other key targets.
Table 3: MAE (↓) of force prediction on the MD17 dataset (kcal/mol/Å). Best results are bolded.
Molecule Aspirin Benzene Ethanol Malonaldehyde
Coord 0.23299 0.15052 0.10909 0.16218
GeoRecon (Ours) 0.21097 0.14734 0.09771 0.15705
[…]
Molecule Naphthalene Salicylic Acid Toluene Uracil
Coord 0.06266 0.13336 0.06843 0.09109
GeoRecon (Ours) 0.05755 0.11996 0.06064 0.08716
[…]
5.2.2 Evaluation on MD17
MD17 benchmark (Chmiela et al., 2017) consists of ab initio molecular dynamics (MD) trajectories for small molecules (e.g., ethanol, malonaldehyde, glycine) at $500~\mathrm{K}$, providing tens of thousands of DFT-computed total energies and atomic forces per molecule. It has become a widely used benchmark for assessing machine-learned force fields in the low-dimensional, gas-phase regime.
As shown in Table 3, GeoRecon consistently improves upon the Coord baseline across all MD17 force prediction tasks. The results demonstrate that incorporating graph-level reconstruction yields clear gains over coordinate denoising alone.
5.2.3 Evaluation on MD22
[…]
Molecule DHA Stachyose AT-AT-CG-CG
Metric Energy Force Energy Force Energy Force
Coord 0.15772 0.13891 0.70198 0.52737 0.48079 0.34858
GeoRecon 0.11955 0.13553 0.59882 0.51666 0.47458 0.34080
[…]
Figure 3: MAE ($\downarrow$) loss curves from linear probing experiments on multiple QM9 tasks. A cosine–warmup learning rate schedule is applied, and both models are trained for 14.7k steps.
We selected three molecules from the MD22 dataset: DHA (56 atoms), Stachyose (87 atoms), and AT-AT-CG-CG (118 atoms) to evaluate our method on small-, medium-, and large-scale systems, respectively. As shown in Table 4, GeoRecon outperforms the Coord backbone on both energy and force prediction tasks across all three molecules, demonstrating the effectiveness of our approach beyond small-molecule scenarios.
5.3 Linear Probing Experiments
In previous work (Kumar et al., 2022), it has been shown that linear probing, that is, appending a lightweight classification head to a frozen, well-pretrained representation model, can achieve stronger performance under out-of-distribution (OOD) settings. Motivated by this finding, we conduct linear probing experiments on GeoRecon and Coord to directly evaluate the influence of different pretraining methods. Since the encoder is kept frozen during these experiments, the degrees of freedom for training are significantly reduced. Consequently, the simplicity of the linear head imposes a strong requirement on the encoder to have already learned sufficiently good features during pretraining.
The experimental results are shown in Figure 3. GeoRecon consistently outperforms Coord across all tasks under the linear probing setting. Notably, Coord exhibits almost no optimization in this setting, suggesting that it can hardly adapt to downstream tasks without full fine-tuning that alters the pretrained structure. As discussed earlier, we attribute this to the limited quality of its pretrained representations. This observation is consistent with the Lipschitz constant measurements reported in the Introduction, further indicating that GeoRecon learns smoother and more informative representations.
6 Conclution and Limitation
Our work presents GeoRecon, a reconstruction-based pretraining framework that bridges node-level denoising and graph-level supervision for 3D molecular representation learning. By conditioning geometry reconstruction on global molecular embeddings, GeoRecon compels the encoder to capture both fine-grained atomic environments and coherent whole-molecule topology. Our theoretical analysis shows that this design produces substantially smoother latent manifolds—reducing local Lipschitz constants by nearly three orders of magnitude compared with node-centric baselines—which provably tightens generalization bounds and improves robustness to structural noise.
Comprehensive experiments across QM9, MD17, and MD22 validate the effectiveness of GeoRecon. The framework achieves competitive results on small-molecule property prediction and consistently improves upon the direct baseline Coord on molecular dynamics prediction tasks (MD17 and MD22) as well as on most targets in QM9. Ablation studies further show that reconstruction noise scaling and decoder depth govern the trade-off between local fidelity and global coherence.
Key advantages of GeoRecon include eliminating dependence on 2D graphs or external labels, seamless integration with SE(3)-equivariant architectures, and a principled connection between representation smoothness and downstream generalization. Looking forward, extending GeoRecon to biomacromolecules, contrastive or generative graph-level objectives, and broader biochemical benchmarks will further test its versatility. Taken together, these results establish geometry-aware graph-level reconstruction as a general paradigm for molecular pretraining, offering a simple yet powerful path toward more sample-efficient and physically grounded representation learning.
Appendix A Ablation Studies
A.1 Ablation of Task Rec and Clean